OpenStudy (anonymous):

multiply, then simplify where possible. 4(3 multiplied by the square root of 2)

OpenStudy (anonymous):

\[12\sqrt{2}\]?

OpenStudy (anonymous):

please show step by step

OpenStudy (anonymous):

u cant. u just multiply 4 by 3

OpenStudy (anonymous):

u dont have to show work for it

OpenStudy (lalaly):

\[4 \times (3\sqrt2)= 4 \times 3 \times \sqrt2 = 12 \sqrt2\]